OpenStudy (anonymous):

Convert 2x^2+9x-6/x^2+x-6 into partial fractions.

OpenStudy (anonymous):

I just need step-by-step guidance, not the answer.

ganeshie8 (ganeshie8):

first divide

ganeshie8 (ganeshie8):

did long division recently ?

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

good. divide so that, degree of numerator becomes less than degree of denominator

OpenStudy (anonymous):

I got 2 + 7x + 6/ x^2 + x -6.

OpenStudy (anonymous):

To get a proper fraction?

OpenStudy (anonymous):

2 + \(\dfrac{7x + 6}{ x^2 + x - 6}\)

ganeshie8 (ganeshie8):

that looks right, next u need to wry about converting : 7x+ 6/ x^2 + x -6. to partial fractions

OpenStudy (anonymous):

(x + 3)(x - 2)

ganeshie8 (ganeshie8):

yes, next setup the fractions

OpenStudy (anonymous):

Do you leave out the 2 that's what I'm confused about...

ganeshie8 (ganeshie8):

yup leave out the 2

ganeshie8 (ganeshie8):

can u wish anything simpler than 2 to integrate ? :) leave 2 as it is, and convert the remaining fraction to partial fractions...

OpenStudy (anonymous):

Okay, what i got \( 2+ \dfrac{4}{x-2} + \dfrac{3}{x+3}\).

OpenStudy (anonymous):

Is that correct?

ganeshie8 (ganeshie8):

how did u get that ?

OpenStudy (anonymous):

Is it wrong?

ganeshie8 (ganeshie8):

it is correct, im asking how did u arrive at that...

OpenStudy (anonymous):

\( 7x + 6=\dfrac{A}{x-2} + \dfrac{B}{x + 3}\)

ganeshie8 (ganeshie8):

wat happened to denominator on left side ?

OpenStudy (anonymous):

What denominator?

ganeshie8 (ganeshie8):

\(\large \frac{7x + 6}{\color{red}{(x+2)(x+3)}}=\dfrac{A}{x-2} + \dfrac{B}{x + 3}\)

ganeshie8 (ganeshie8):

^^that denominator

OpenStudy (anonymous):

I didn't write it.

OpenStudy (anonymous):

I left it by mistake.

ganeshie8 (ganeshie8):

ok then fine :)

ganeshie8 (ganeshie8):

i see ur doing it right... :)

OpenStudy (anonymous):

I didn't get it when i posted it up but then got it after.

OpenStudy (anonymous):

Thanks :)

ganeshie8 (ganeshie8):

good :P